If 𝐶̅ = R25 billion, I = R35 billion and c = 0.80


The marginal propensity to save is …

[1] 0.80

[2] 1

[3] 1.20

[4] 0.20


The value of the multiplier is …

[1] 5

[2] 0.20

[3] 1

[4] 2


The equilibrium level of income is …

[1] 60

[2] 120

[3] 300

[4] 90


1
Expert's answer
2021-06-24T12:04:47-0400

Solution:

a.). The marginal propensity to save is 4.) 0.20.

c = MPC = 0.8

MPS = 1 – MPC = 1 – 0.8 = 0.2


b.). The value of the multiplier is 1.). 5.

Multiplier = "\frac{1}{(1 - MPC)} = \frac{1}{(1 - 0.8)} = \frac{1}{0.2} = 5"

 

c.). The equilibrium level of income is 3.). 300.

At equilibrium: AD = AS

Y = C + I + G

C = 25 + 0.80Y

I = 35

Y = C + I

Y = 25 + 0.80Y + 35

Y – 0.8Y = 25 + 35

0.2Y = 60

Y = "\frac{60}{0.2} = 300"


The equilibrium level of income (Y) = 300 billion.